# Shapeng Jiang Lijia Wei Chen Zhang

### v1

@@ 3.2 Variance-covariance of responses and similarity of correlations
In Figure 2, most cells are shaded light grey, indicating complete agreement, with only two instances of complete disagreement in the China sample regarding social values. In the U.S. sample, only 3 out of 51 pairwise correlations indicate partial disagreement, while in the China sample, 5 out of 51 correlations show partial disagreement. This reflects that the LLM-generated responses effectively preserve the correlations between various human beliefs and attitudes.
While synthetic data may seem fairly accurate in aggregate, some issues are still evident when examining conditional relationships. To assess whether the correlational structure of the synthetic data aligns with that of the human benchmark, we compare regression outcomes using actual and synthetic responses as the dependent variable.
$\text{Response}_{i,j}=\beta_{0}+D_{j}\mathbf{\beta_{1}}\mathbf{X}_{i}+(1-D_{j})\mathbf{\beta_{2}}\mathbf{X}_{i}+\epsilon_{i,j}$ (1)
where $i$ represents respondents, $j$ denotes the data source (human or LLM sample). $D_{j}$ is a dummy variable that equals 1 for humans and 0 for LLMs. $\mathbf{X}_{i}$ is a vector containing respondent characteristics used in the persona prompt (age, gender, education, income, marital status, occupation). Our primary focus is on the coefficient vectors $\mathbf{\beta_{1}}$ and $\mathbf{\beta_{2}}$, which capture the partial correlations between each covariate in $\mathbf{X}_{i}$ and the response for human and LLM samples, respectively.
We run the specification in Equation 1 for each response (questions on social values and trust), totaling 15 regressions. Figure 3 displays the resulting coefficient estimates, with human coefficient, $\mathbf{\beta_{1}}$, on the x-axis and LLM coefficient, $\mathbf{\beta_{2}}$, on the y-axis, organized by independent variables. Points closer to the 45-degree line indicate an alignment in partial correlation observed from either human or synthetic data. In the off-diagonal quadrants (upper left and bottom right), points reflect coefficients with reversed signs depending on the data source, indicating that conclusions based on human data would be opposite to those drawn from the LLM sample.
Overall, most coefficients fall close to the 45-degree line, with the exception of dummy variables related to industry occupation—agricultural and industrial workers—in panels (h) and (i). This indicates that the LLM is fairly effective in predicting an individual’s responses based on these characteristics. These findings support the feasibility of using demographic features to predict other individual behaviors, laying a solid foundation for our subsequent prediction task.
(a) Age
(b) Gender
(c) Secondary education
(d) Tertiary education
(e) Unpartnered
(f) Partnered
(g) Income level
(h) Agricultural worker
(i) Industrial worker
(j) Service worker
Figure 3: Within-subject correlations across LLM and human.
3.3 Predicting the outcome of 2024 U.S. presidential election
Using the prompts from parts (b) and (c) in Appendix A, we enabled the LLM to simulate the voting behavior of each ANES respondent. Based on different prompts and calculation methods, we obtained three measures of predicted results. The first two use the role-play prompt and the structural prompt, respectively. Although the historical voting prompt directs the LLM to leverage historical data, how the model utilizes its vast stored information remains a black box to us. Therefore, the third method combines the predictions from the role-play prompt with the actual historical vote shares, weighting them to create a more balanced forecast. Specifically, the predicted vote share for party $p$ in state $i$ for year $y$ is defined as follows:
@@
where $h$ denotes the weight. We fine-tuned the weights until the predicted outcomes closely aligned with the actual state-level winners. We then applied this optimized weighting parameter $\hat{h}$ to forecast the outcome of the 2024 presidential election. All the results in Section 3.3 was obtained on or before October 30, 2024, approximately one week before Election Day on November 5, 2024.
In Figure 4, panel (a) presents the results using the role-play prompt, while panel (b) shows the results using the structural prompt. In panel (c), $\hat{h}$ was set to 0.88 to align the predictions more closely with the actual election results in 2016. In panel (d), we calculated the 2016 election outcome within the ANES sample based on respondents’ self-reported voting behavior. This approach allows us to estimate how the election would have played out using the ANES data, offering a benchmark for comparison. Due to sampling errors, the survey benchmark do not perfectly match the actual election outcomes. We placed the actual election results in panel (e) for comparison purposes, and found that the ANES sample results in 2016 differed from the actual election outcome in 8 states. Among all the predictions, the structural prompt performed the best, with Michigan and Florida being the only states it mispredicted. Although the role-play prompt showed weaker predictive performance, incorrectly forecasting 8 states, we improved its accuracy by incorporating historical voting data. With this adjustment, it closely approximated the actual outcome, matching the structural prompt by incorrectly predicting the same number of states: Wisconsin and Michigan.
(a) Role-play prompt
Dem (343) vs Rep (195)
(b) Structural prompt
Dem (278) vs Rep (260)
(c) Role-play prompt with adjustment ($\hat{h}=0.88$)
Dem (259) vs Rep (279)
(d) ANES 2016
Dem (297) vs Rep (241)
(e) 2016 actual voting results
Dem (227) vs Rep (304)
In Figure 5, the structural prompt also performed the best, with Arizona as the only incorrect prediction. The role-play prompt, however, mispredicted 8 states. After adjustment, it closely aligned with the actual results, missing just one more state (North Carolina) compared to the structural prompt. Both Arizona and North Carolina are swing states. Surprisingly, the 2020 ANES post-election survey performed poorly, mispredicting the outcomes in 8 states.
(a) Role-play prompt
Dem (406) vs Rep (132)
(b) Structural prompt
Dem (295) vs Rep (243)
(c) Role-play prompt with adjustment ($\hat{h}=0.8$)
Dem (310) vs Rep (228)
(d) ANES 2020
Dem (333) vs Rep (205)
(e) 2020 actual voting results
Dem (306) vs Rep (232)
Finally, using data from 6,571 respondents from the ANES 2020 survey, we predicted the outcome of the 2024 election. All three measurement methods consistently forecast that Donald Trump would win the election, and the Republican would receive approximately 300 of the 538 total electoral votes. Detailed state-by-state Democratic and Republican vote shares are available in Table LABEL:tab:predict_2024 in the Appendix B. Columns (5) and (6) include polling data from the website 270toWin [^2] See https://www.270towin.com/..
(a) Role-play prompt
Dem (220) vs Rep (318)
(b) Structural prompt
Dem (252) vs Rep (286)
(c) Role-play prompt with adjustment ($\hat{h}=0.8$)
Dem (229) vs Rep (309)
4 Discussion and Implications
This study demonstrates that large language models exhibit significant in-sample predictive capabilities, closely mirroring human responses when demographic variables are input. For out-of-sample predictions, such as elections, the model also shows potential in forecasting trends. These findings suggest that LLMs could serve as effective supplements to traditional survey methodologies. First, by accurately simulating human responses based on demographic inputs, LLMs can reduce survey costs, serving as a supplementary data source when direct collection is limited. They also help extend survey reach, filling in data gaps for underrepresented groups. Additionally, LLMs offer a testing ground to validate and refine survey questions, ensuring relevance to target populations. Lastly, their predictive capacity allows researchers to anticipate trends, especially in evolving social issues.
While this study demonstrates the potential of LLMs in simulating human responses, there remain areas for further exploration. Our analysis primarily used ChatGPT-4o, without assessing other LLMs, which may exhibit different biases on value-laden topics due to varied training data. Additionally, our data is drawn from the WVS and ANES surveys, so incorporating more diverse sources could enhance generalizability. Lastly, the cross-cultural analysis is centered on the U.S. and China; future studies could broaden this scope to better understand LLM performance across diverse cultural contexts.
Appendix A Prompts
(a) Simulation prompt
@@ Appendix A Prompts
    “For the following questions, choose 1 for Agree strongly, 2 for Agree, 3 for Neither agree nor disagree, 4 for Disagree, and 5 for Disagree strongly. How would you feel about the following statements? Do you agree or disagree with them?”
    • Q33 When jobs are scarce, men should have more right to a job than women
    • Q34 When jobs are scarce, employers should give priority to people of this country over immigrants
    • Q35 If a woman earns more money than her husband, it’s almost certain to cause problems
    • Q36 Homosexual couples are as good parents as other couples
    • Q37 It is a duty towards society to have children
    • Q38 Adult children have the duty to provide long-term care for their parents
    • Q39 People who don’t work turn lazy
    • Q40 Work is a duty towards society
    • Q41 Work should always come first, even if it means less spare time
@@
    “For the following questions, choose 1 for Trust completely, 2 for Trust somewhat, 3 for Do not trust very much, 4 for Do not trust at all. I’d like to ask you how much you trust people from various groups. Could you tell me for each whether you trust people from this group completely, somewhat, not very much or not at all?”
    • Q58 Your family
    • Q59 Your neighborhood
    • Q60 People you know personally
    • Q61 People you meet for the first time
    • Q62 People of another religion
    • Q63 People of another nationality
@@
    • Q91 Five countries have permanent seats on the Security Council of the United Nations. Which one of the following is not a member?
@@
    • Q92 Where are the headquarters of the International Monetary Fund (IMF) located?
@@
    Ethical norms and values
@@
    • Q178 Avoiding a fare on public transport
    • Q179 Stealing property
    • Q180 Cheating on taxes if you have a chance
    • Q181 Someone accepting a bribe in the course of their duties
    • Q182 Homosexuality
    • Q185 Divorce
    • Q190 Parents beating children
    • Q192 Terrorism as a political, ideological or religious mean

### v2

@@ Results
As shown in Figure 3, in the U.S. sample, only 3 out of 51 pairwise correlations indicate partial disagreement. In contrast, in the China sample, 22 out of 51 correlations show partial disagreement, while 2 exhibit complete disagreement. This reflects that the LLM-generated responses effectively preserve the correlations between various human beliefs and attitudes in the U.S. context, whereas nearly half of the correlations are not retained in China. After applying the Matching-LLM method, the correlations in the U.S. context remain consistent. Although only a subset of questions can be tested due to changes in survey items between WVS6 and WVS7, we find notable improvements within this subset in the China sample. Overall, in both the U.S. and China samples, the majority of circles fall in the first and third quadrants, indicating that the correlations between responses in the human sample are preserved by both the LLM and Matching-LLM approaches.
WVS survey responses: out-of-sample prediction
We find that using LLM alone yields reliable prediction results, and incorporating historical data for weighting further enhances the model’s ability to replicate human responses. However, thus far, the weights used in the Matching-LLM approach were optimized within this sample using a least-squares method. Specifically, the weights in the Matching-LLM approach were derived by minimizing the Euclidean distance between the Matching-LLM response vector and the human response vector. The response vector includes all the survey questions in Figure 1. Therefore, for individual questions, the Matching-LLM approach may not necessarily produce results closer to human responses than the raw LLM-generated responses. However, from the perspective of all survey questions as a whole, the Matching-LLM approach is a convex combination of LLM-generated data and human data, making its superior performance over the raw LLM approach a certainty.
To further assess the ability of the Matching-LLM approach to simulate human responses, we tested its performance on new survey questions by using the weights derived in Figure 1. In other words, we treated the survey responses simulated thus far as the training set to optimize the weights and used previously unexamined WVS survey responses as the test set to evaluate the performance of the Matching-LLM approach.
Figure 4 compares human responses with LLM-generated responses and Matching-LLM responses using the weights obtained from the training set and new WVS questions related to gender roles and family values (Method). We find that, except for the “work-motherhood impact” question in the U.S. sample, the responses generated by the Matching-LLM approach are consistently closer to human responses than those generated by the raw LLM approach. This finding reinforces confidence in the Matching-LLM approach.
Figure 4: Out-of-sample prediction: WVS survey responses.
However, these evaluations are still conducted within the WVS sample, relying on existing data to examine the predictive capabilities of the LLM and Matching-LLM approaches. In what follows, we extend our analysis to out-of-sample prediction by leveraging an upcoming real-world event—the 2024 U.S. presidential election—to test the predictive performance of both approaches.
2016 and 2020 U.S. election: in-sample prediction
We began by using data from the ANES surveys to replicate the outcomes of the 2016 and 2020 U.S. presidential elections. The LLM was instructed to adopt the persona of each ANES respondent and simulate their vote for one of the two major party candidates in the respective election. The ballots were designed to closely resemble the format used in the corresponding U.S. presidential election of that year. The individual-level results were then aggregated at the state level to determine the winner in each state, and the electoral votes were summed across states to identify the overall winner.
When comparing the LLM-generated results to the self-reported voting outcomes from ANES respondents, this constitutes an in-sample prediction. Our focus, however, is on out-of-sample prediction—comparing the LLM-generated state-level results to the actual voting outcomes. Additionally, we applied a weighting method using historical data to improve the LLM’s predictive accuracy. Specifically, we weighted the state-level aggregated results from the LLM prediction with historical voting outcomes for each state. The weights were fine-tuned until the Matching-LLM approach aligned as closely as possible with state-level outcomes from past elections.
In Figure 5, panel (a) placed the actual election results for comparison purposes, while panel (b) shows the 2016 election outcome within the ANES sample based on respondents’ self-reported voting choices, which indicates how the election would have played out using the ANES data, offering a benchmark for comparison. In panel (c), we directly prompted LLM to take the personas and casted their votes. Panel (d) displays the results generated by Matching-LLM approach, where the weight was calibrated to closely align predictions with the actual 2016 election results.
Due to sampling errors, the survey benchmark in panel (b) do not necessarily match the actual election outcomes. Compared with panel (a), we found that the ANES sample results differed from the actual 2016 election outcome in 8 states. Similar to ANES sample, the LLM approach incorrectly forecasted 8 states. We improved the LLM’s accuracy by incorporating historical voting data in panel (d), the Matching-LLM approach closely approximated the actual outcome, incorrectly predicting only two states: Wisconsin and Michigan.
Figure 6 reveals a pattern similar to the replication of the 2016 election. In panel (b), the 2020 ANES post-election survey underperformed, incorrectly predicting outcomes in 8 states. Similarly, the LLM prompt alone also mispredicted 8 states. However, after adjustment, the Matching-LLM results closely matched the actual outcomes, missing only two states: Arizona and North Carolina, both of which are swing states.
Given that Donald Trump participated in the last three U.S. presidential elections, the historical data used in the Matching-LLM approach comes from the actual results of the 2016 and 2020 elections. As a result, regardless of the chosen weight, Matching-LLM inevitably produces more accurate predictions compared to the LLM alone. Therefore, we must carefully interpret the rationale for using the Matching-LLM approach in predicting the 2016 and 2020 elections.
2024 U.S. election and 2025 German federal election: out-of-sample prediction
In this section, the purpose of employing the Matching-LLM approach is not merely to replicate known election outcomes but to derive an appropriate historical weighting scheme for predicting the 2024 U.S. presidential election and the 2025 German federal election.
To predict the 2024 election, we leveraged the demographic characteristics of respondents from ANES 2020. We derived the weights for the Matching-LLM method from in-sample election prediction, and it was then applied to forecast the outcome of the 2024 U.S. presidential election. Notably, all results in this section were obtained on or before October 30, 2024, approximately one week before Election Day on November 5, 2024. Furthermore, the corresponding working paper draft was publicly available on SSRN [^4] See https://ssrn.com/abstract=5008330. prior to the election, ensuring the predictions were made entirely ex-ante, without any retrospective adjustments based on actual results.
Using data from 6,571 respondents from the ANES 2020 survey, we predicted the outcome of the 2024 election. Both LLM and Matching-LLM methods consistently forecast that Donald Trump would win the election. Detailed state-by-state Democratic and Republican vote shares are available in Table LABEL:tab:predict_2024 in the Appendix B. Columns (5) and (6) include polling data from the website 270toWin [^5] See https://www.270towin.com/..
In terms of electoral votes, both approaches produced predictions that closely matched the final outcome, with the Matching-LLM approach demonstrating superior accuracy. Donald Trump, representing the Republican Party, secured 312 electoral votes, just three votes more than the Matching-LLM prediction of 309. The standard LLM method mispredicted six states, whereas the Matching-LLM approach exhibited remarkable predictive precision, making errors in only two states: Nevada and New Hampshire. Notably, among the seven key swing states identified as crucial before the election (Arizona, Georgia, Michigan, Nevada, North Carolina, Pennsylvania, and Wisconsin), the Matching-LLM approach mispredicted only Nevada, further underscoring its effectiveness in capturing electoral dynamics.
In the forecasting of 2025 German federal election, we employed individual-level data from the CESIS database666www.gesis.org—a comprehensive repository that captures detailed demographic, socioeconomic, and political information from German voters, comprising responses from 3,965 individuals—to adjust GPT-based vote share simulations for the upcoming 2025 German federal election. By comparing the GPT-simulated results with actual historical outcomes from 2017 and 2021, we derived weighting values for each party and applied these to the 2025 predictions. After weighting the adjusted figures, our forecasts indicate that the SPD is projected to receive approximately 19.55% of the vote, while the CDU/CSU is expected to lead with about 33.11%. The Greens are forecast at 15.17%, and the FDP at 3.77%. Notably, the AfD is predicted to secure around 23.89%, and Die Linke about 4.51%. These weighted results demonstrate that our methodology effectively compensates for the systematic biases observed in the raw GPT simulations, yielding a forecast that more accurately reflects historical voting trends and provides a robust prediction for the 2025 German election.
Discussion
This study demonstrates that the latest LLM exhibit significant predictive capabilities, closely mirroring human responses when demographic variables are input. For out-of-sample predictions, the model also show potential in prediction the survey responses and US election outcomes. Additionally, based on the dual-process theory of fast and slow thinking, this study develops a method named Matching-LLM that combines historical data with synthetic data generated by the LLM to improve their predictive capabilities. Results show the Matching-LLM approach outperforms the use of LLMs alone in terms of prediction accuracy.
These findings suggest that LLMs could serve as effective supplements to traditional survey methodologies (Jansen et al.,, 2023; Rossi et al.,, 2024). By accurately predicting human responses based on demographic inputs, LLMs can reduce survey costs and serve as a supplementary data source when direct collection is limited. They also help extend survey reach, filling in data gaps for underrepresented groups. Furthermore, LLMs offer a testing ground to validate and refine survey questions, ensuring relevance to target populations. Lastly, their predictive capacity allows researchers to anticipate trends, especially in evolving social issues.
This study extends prior research on using LLMs to simulate survey responses (Bisbee et al.,, 2023; Al Tamime et al.,, 2024; Kim et al.,, 2024). First, we focused on a set of representative large-scale public opinion questions, such as work ethic, gender roles, family values, and trust. This approach is more comprehensive than similar studies and enhances confidence in using LLMs to assist large-scale survey research. Additionally, unlike prior research that focused solely on the accuracy of LLM predictions, we evaluated whether LLMs can replicate relationships between variables and across samples. We also engage with election prediction studies. Traditional election forecasts rely on macro features such as economic indicators, national opinion polls, and social media data, often analyzed using machine learning or regression models (Kennedy et al.,, 2017; Brito et al.,, 2021). For this study, limited computational resources led us to use survey data as a demographic source. However, LLMs could theoretically leverage national census data to minimize sampling errors. Moreover, ChatGPT-4o, trained on vast amounts of online data up to October 2024, offers unparalleled access to data compared to traditional methods. The primary limitation of using LLMs in election prediction, however, lies in the “black box” nature of their prediction process, which obscures the specific factors influencing voter choices.
The Matching-LLM approach proposed in this study offers a novel perspective on LLM application. While LLMs are generally recognized for effectively predicting human language and reasoning, their training on vast textual data introduces a bias toward rational thinking, limiting their ability to capture intuitive, non-rational human behavior. This motivated the development of the Matching-LLM approach, which creates a convex combination of historical data and LLM-generated data. Since human behavior is governed by both fast (intuitive) and slow (deliberative) thinking systems, this combined approach better mimics real-world human behavior than either system alone. However, this innovative approach carries risks. When mirroring human behavior, biases from slow thinking arise not only from its rational nature but also from adherence to ethical norms, laws, and progressive ideologies(Feng et al.,, 2023; Scherrer et al.,, 2024). Our results show that most deviations between LLM-simulated and human responses occur on such value-laden topics. For example, LLMs tend to be more moralistic, law-abiding, and progressive on issues like LGBTQ+ rights and gender equality. Conversely, biases in fast thinking often stem from self-reporting errors driven by social desirability bias(Phillips and Clancy,, 1972), such as income exaggeration. For instance, in the 2024 U.S. presidential election, discrepancies between opinion polls and the final results—where many polls predicted Harris’s victory—might have been due to sampling biases or reluctance among Trump supporters to participate in surveys (Boyle et al.,, 2023; Graefe,, 2024). When both thinking systems produce errors, the direction of these errors determines the efficacy of the convex combination. Opposing biases may cancel each other out, while aligned biases could amplify inaccuracies.
While this study demonstrates the potential of LLMs in predicting human responses, areas for further exploration remain. Our analysis primarily used ChatGPT-4o, without examining other LLMs that may exhibit different biases on value-laden topics due to varied training data. Moreover, our data relied on WVS and ANES surveys, and incorporating more diverse sources could improve generalizability. Lastly, our cross-cultural analysis was centered on the U.S. and China. Future studies should expand this scope to evaluate LLM performance across a broader range of cultural contexts.
Methods
We confirm that our research adheres to all applicable ethical guidelines. As the datasets used were derived from publicly available sources, participant compensation does not apply to this study. Additionally, none of the studies involved pre-registration.
Datasets
The LLM model, ChatGPT-4o, was trained on data up to October 2023, and we relied exclusively on the ChatGPT API tool, without incorporating any real-time search capabilities. In the survey task, when using LLM approach, we provided ChatGPT with demographic characteristics of respondents from the United States ($N=2077$) and China ($N=2001$), as recorded in the 7th Wave of the World Values Survey (WVS). After adopting a persona defined by a set of demographic characteristics of a respondent, the LLM was tasked with generating a dataset of synthetic opinions on social values, trust, common-sense questions, and ethical norms and values.
In the election task, when using LLM approach, the LLM was provided with the demographic characteristics of respondents from ANES2020 ($N=6571$). We selected ANES2020 for three main reasons: first, its sample siz is about twice that of ANES2016, making it more representative; second, it includes information on respondents’ voting choices in the 2016 election; and third, as of the 2024 U.S. presidential election, ANES2024 had only released a pilot study with approximately 1,900 respondents, which was insufficient for analysis.
The LLM simulated voting behavior by adopting the persona of an ANES respondent and casting votes in three elections: Hillary Clinton vs. Donald Trump (2016), Joe Biden vs. Donald Trump (2020), and Kamala Harris vs. Donald Trump (2024). Following U.S. election rules, the simulated votes were aggregated by state using the survey’s sampling weights to ensure representativeness. These state-level results were then combined nationally to calculate electoral votes and determine the winner. This process allows us to compare the LLM-generated synthetic voting results with the actual outcomes of the elections.
Matching and weights.
When employing the Matching-LLM approach, we first performed matching and then determined a unified weight for combining historical data with LLM-generated data. The methods for matching and weighting differ between predicting WVS survey responses and U.S. election voting behavior.
The Matching-LLM method consists of three steps:
  1. 1.
    Matching: This step involves using historically collected individual characteristic data to match with the features of current respondents. The goal is to generate a corresponding “System 1” response for each respondent. If the data structure follows a panel format, the historical response data is directly matched.
  2. 2.
    LLM: In this step, a large language model is employed to generate the “System 2” response for each respondent. All simulated responses generated by the LLM were obtained via the API rather than the standard visual interface.
  3. 3.
    Weighting: The final step involves weighting the responses from System 1 and System 2. For in-sample predictions, the weights are optimized to achieve the best possible results. For out-of-sample predictions, the weighting is determined by referencing the weights used for similar question types in the in-sample setting. The magnitude and application of these weightings depend on the data structure and the specific research question. For instance, when working with panel data, historical data are typically assigned a higher weight because they provide a richer and more consistent set of relevant information. In contrast, for non-panel data, the weighting applied to historical data tends to be significantly lower. Additionally, it is essential to consider the potential for polarization in ChatGPT’s responses on certain topics.
As a example, for the prediction of WVS survey responses, we begin with LLM-synthesized responses generated based on WVS7 individual characteristics. Using propensity score matching (PSM), we identify the most similar individual from WVS6 for each individual in WVS7 through 1:1 nearest-neighbor matching. The demographic variables used for matching include age, gender, education level, marital status, occupation, and income scale.
We then combine the LLM-synthesized responses from WVS7 and the responses of matched individuals from WVS6 using the following equation:
$R_{k}^{\text{Matching-LLM}}=hR_{k}^{\text{WVS6}}+(1-h)R_{k}^{\text{WVS7-LLM}}$
where $R_{k}^{\text{Matching-LLM}}$ is the response vector generated using the Matching-LLM approach for the $k$-th question, $R_{k}^{\text{WVS6}}$ represents the response vector from individuals matched in WVS6, and $R^{\text{WVS7-LLM}}$ contains the LLM-synthesized responses from WVS7. $h\in[0,1]$ denotes the weight.
We employ the least squared method to estimate the weight $\hat{h}$:
$\hat{h}=argmin\frac{1}{K}\sum_{k}\sqrt{(R_{k}^{\text{Matching-LLM}}-R_{k}^{\text{WVS7}})^{T}\cdot(R_{k}^{\text{Matching-LLM}}-R_{k}^{\text{WVS7}})}$
such that the average Euclidean distance between the response vectors $R_{k}^{\text{Matching-LLM}}$ and $R_{k}^{\text{WVS7}}$ is minimized, where $R_{k}^{\text{WVS7}}$ denotes the actual responses from WVS7. We conducted the matching and weight generation processes separately for the China and U.S. samples, resulting in weights $\hat{h}=0.31$ for the China sample and $\hat{h}=0.23$ for the U.S. sample.
For the prediction of U.S. election using the ANES2020 sample, since our focus is on state-level prediction accuracy, we simply weight the actual vote shares of the two-party candidates at the state level (from the 2016 and 2020 elections) and the LLM-synthesized vote shares. Specifically, the predicted vote share using Matching-LLM approach for party $p$ in state $i$ for year $y$ is defined as follows:
@@ Methods
where $h$ denotes the weight. We fine-tuned $h$ until the predicted outcomes aligned as closely as possible with the actual state-level winners. We then applied this optimized weighting parameter $\hat{h}$ to forecast the outcome of the 2024 presidential election.
GPT-4o prompt
The prompt used in the LLM approach is shown below, with the demographics used detailed in the appendix.
LLM prompt: WVS
@@
    “How would you feel about the following statements? Do you agree or disagree with them? Choose 1 for Agree strongly, 2 for Agree, 3 for Neither agree nor disagree, 4 for Disagree, and 5 for Disagree strongly. ”
    • Male-preferred employment. When jobs are scarce, men should have more right to a job than women
    • Domestic worker priority. When jobs are scarce, employers should give priority to people of this country over immigrants
    • Male as breadwinner. If a woman earns more money than her husband, it’s almost certain to cause problems
    • Homosexual parents. Homosexual couples are as good parents as other couples
    • Parenthood as social duty. It is a duty towards society to have children
    • Filial duty. Adult children have the duty to provide long-term care for their parents
    • Anti-laziness. People who don’t work turn lazy
    • Work as social duty. Work is a duty towards society
    • Work over leisure. Work should always come first, even if it means less spare time
@@
    “I’d like to ask you how much you trust people from various groups. Could you tell me for each whether you trust people from this group completely, somewhat, not very much or not at all? Choose 1 for Trust completely, 2 for Trust somewhat, 3 for Do not trust very much, 4 for Do not trust at all. ”
    • Trust on family. Your family
    • Trust on neighbors. Your neighborhood
    • Trust on acquaintances. People you know personally
    • Trust on strangers. People you meet for the first time
    • Trust on religious outgroups. People of another religion
    • Trust on foreigners. People of another nationality
@@
    • Question of the UN. Five countries have permanent seats on the Security Council of the United Nations. Which one of the following is not a member?
@@
    • Question of the IMF. Where are the headquarters of the International Monetary Fund (IMF) located?
@@
    Ethical norms and laws
@@
    • Fare evasion acceptance. Avoiding a fare on public transport
    • Theft acceptance. Stealing property
    • Tax evasion acceptance. Cheating on taxes if you have a chance
    • Bribery acceptance. Someone accepting a bribe in the course of their duties
    • Homosexuality acceptance. Homosexuality
    • Divorce acceptance. Divorce
    • Parental abuse acceptance. Parents beating children
    • Terrorism acceptance. Terrorism as a political, ideological or religious mean
  5. 5.
    Out-of-sample questions in Figure 4
    “For each of the following statements, can you tell me how strongly you agree or disagree with each. Do you strongly agree, agree, disagree, or strongly disagree? Choose 1 for Strongly agree, 2 for Agree, 3 for Disagree, and 4 for Strongly disagree.”
    • Parental pride goal. One of my main goals in life has been to make my parents proud
    • Work-motherhood impact. When a mother works for pay, the children suffer
    • Male political leadership. On the whole, men make better political leaders than women do
    • University gender gap. A university education is more important for a boy than for a girl
    • Male business executives. On the whole, men make better business executives than women do
    • Housewife fulfillment. Being a housewife is just as fulfilling as working for pay
LLM prompt: ANES
Step 1: Assigning a persona to the LLM
Step 2: Setting the scenario
“Hello!
U.S. Presidential and Vice Presidential Election - Vote for One Party
Step 3: Presenting options and gathering responses
  1. 1.
    KAMALA D. HARRIS / TIM WALZ (Democratic)
  2. 2.
    DONALD J. TRUMP / J.D. VANCE (Republican)
  • 2016 United States presidential election
    1. 1.
      HILLARY R. CLINTON / TIMOTHY M. KAINE (Democratic)
    2. 2.
      DONALD J. TRUMP / MICHAEL R. PENCE (Republican)
  • 2020 United States presidential election
    1. 1.
      JOSEPH R. BIDEN / KAMALA D. HARRIS (Democratic)
    2. 2.
      DONALD J. TRUMP / MICHAEL R. PENCE (Republican)
Appendix A Demographics